Question 1: Solve the following equation:

9 + 3 e 2 x + 1 = 12 − e 2 x + 1

A. approximately x = 1.305

B. approximately x = -0.644

C. approximately x = 1.674

D. approximately x = -0.562

E. none of the above


choose the correct answer



1
Expert's answer
2020-08-24T07:52:45-0400
"9+3e^{2x+1}=12-e^{2x+1}"


"e^{2x+1}=0.75"

"2x+1=ln0.75"


"x=0.5(ln0.75-1)"


"x=-0.643841"

Answer B